Ellipsis, Underspecification, Events and More in Dynamic

Semantics

Jeroen Groenendijk (editor)

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852 Deliverable R2.2.C September 1995

Ellipsis, Underspecification, Events and More in Dynamic

Semantics

Jeroen Groenendijk Universiteit van Amsterdam

(editor)

authors: Helle Frisak Sem Reinhard Muskens

Tim Fernando Regine Eckardt Fritz Hamm

Jeroen Groenendijk, Martin Stokhof and Frank Veltman Paul Dekker

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852 Deliverable R2.2.C September 1995

esprit Basic Research Project 6852 dyana-2

Dynamic Interpretation of Natural Language

Universiteit van Amsterdam Institute for Logic, Language and Computation (ILLC)

University of Edinburgh Centre for Cognitive Science (ECCS)

Universität München Centrum für Informations- und Sprachforschung (CIS)

Universitetet i Oslo Department of Linguistics and Philosophy (ILF)

Universität Stuttgart Institut für Maschinelle Sprachverarbeitung (IMS)

Universität Tübingen Seminar für Sprachwissenschaft (SfS)

Universiteit Utrecht Research Institute for Language and Speech (OTS)

For copies of reports, updates on project activities and other dyana-related information, contact:

The dyana-2 Project Administrator ILLC/Department of Philosophy University of Amsterdam Nieuwe Doelenstraat 15 NL-1012 CP Amsterdam

http://www.fwi.uva.nl/research/illc/dyana/Home.html

\begin{matrix}c \\ ◯ \end{matrix} 1995, The Individual Authors

No part of this document may be reproduced or transmitted in any form, or by any means, electronic or mechanical, including photocopy, recording, or any information storage and retrieval system, without permission from the copyright owner.

Contents

Introduction by the editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v

Task 2.2, subtask 4: VP-Ellipsis VP Ellipsis beyond Proper Name Subjects Helle Frisak Sem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Task 2.2, subtask 3: Quantificational Scope Ambiguities Order-Independence and Underspecification Reinhard Muskens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

Additional Contribution A Logical Connective for Ambiguity Requiring Disambiguation Tim Fernando . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

Task 2.2, subtask 6: Events and Adverbial Modification Does Information Flow in Event Structures Regine Eckardt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

Additional Contribution Perfect and Imperfect Nominals Fritz Hamm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

Additional Contribution Coreference and Contextually Restricted Quantification Jeroen Groenendijk, Martin Stokhof and Frank Veltman. . . . . . . . . . . . 97

Additional Contribution On Context and Identity Paul Dekker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

Introduction

As the title indicates, this last deliverable within the task Extending the empirical domain of the area Interpretation in context, covers a wide range of topics.

Helle Frisak Sem’s contribution ‘VP-ellipsis beyond proper name subjects’ is a continuation of her work on VP-ellipsis within the framework of DRT reported in Deliverable R2.2.B. She discusses the extension of her earlier analysis to cases which involve scope and quantification. Her investigation leads to the conclusion that an inferential or substitutional approach along the lines of her earlier proposal, which assumes an intermediate representational level, is better suited to deal with these phenomena than the alternative approach which is based on making an identification at the level of predication.

Reinhard Muskens’ paper presents an LFG-like annotated phrase structure formalism which connects underspecified DRT-like semantic representations with other components of the grammar. The underspecified nature of the semantic representations makes it possible to postpone semantic disambiguation to the point where contextual or other information gives sufficient clues as to decide upon a particular reading. A second feature of the proposed framework is that interpretation is order-independent, i.e., that it does not decide beforehand whether meanings are assigned top-down—which is the preferred order from the point of view of efficient parsing—or whether it is bottom-up— which is the preferred order within a compositional semantics. These features of order-independence and underspecification are important to arrive at efficient combined syntactic and semantic processing.

The (additional) contribution ‘A logical connective for ambiguity requiring disambiguation’ by Tim Fernando is also connected to the issue of ambiguity and underspecification. A naive way to represent ambiguity is in terms of an ordinary logical disjunction of the readings in question. Fernando sets out to investigate the logical possibilities and properties of a connective which, unlike ordinary disjunction, could be used in a proper underspecified representation of ambiguous expressions.

Regine Eckardt’s paper ‘Does information flow in event structures’ concerns the subtask on events and adverbial modification. In particular, she adresses temporal and spatial modifications of complex (‘conjunctions’ of) events, and notes that under the existing ontologies of complex events, such modifications cannot be properly dealt with. In order to avoid these problems, Eckardt introduces a more fine-grained and dynamic non-Boolean interpretation of conjunctions of event descriptions.

The ontology of events also plays a role in the (additional) contribution ‘Perfect and imperfect nominals’ by Fritz Hamm. The paper is concerned with the (onto)logical status of two classes of nominalized predicates, Vendler’s perfect and imperfect nominals, and of two kinds of verbal contexts, Vendler’s narrow and loose containers, which discrminate between these two kinds of nominals. Hamm proposes a type-free intensional approach combined with a Davidsonian events semantics in order to arrive at a logical reconstruction of Vendler’s observations.

The last two papers in this deliverable are both additional contributions, and both concern further extensions of the systems of update semantics developed within DYANA.

The paper ‘Coreference and contextually restricted quantification’ by Jeroen Groenendijk, Martin Stokhof and Frank Veltman, presents a descriptive extension of their earlier work on dynamic epistemic predicate logic, reported in deliverable R2.1.B. They show that the framework developed in their earlier papers naturally gives rise to a notion of contextually restricted quantification, which can be used, among other things, in an analysis of anaphoric definite descriptions.

In his paper ‘On context and identity’ Paul Dekker analyzes a number of traditional logical-philosophical problems concerning identity and demonstrative reference within update semantics. He shows that in order to give a proper analysis of these problems it is not sufficient to model information as such, but that it is essential to tie the information of agents to the perceptual properties of the actual situations in which they find themselves. Dekker shows that an update semantics making use of such situated information states provides us with better tools for the analysis of reference and identity.

Amsterdam, September 1995 Jeroen Groenendijk

Task 2.2, subtask 4 VP-Ellipsis

VP Ellipsis beyond Proper Name Subjects

Helle Frisak Sem

(Universitetet i Oslo)

Task 2.2, subtask 3 Quantificational Scope Ambiguities

Order-Independence and Underspecification

Reinhard Muskens

(Katholieke Universiteit Brabant)

Additional Contribution

A Logical Connective for Ambiguity Requiring Disambiguation

Tim Fernando

(Universität Stuttgart)

Task 2.2, subtask 6 Events and Adverbial Modification

Does Information Flow in Event Structures

Regine Eckardt

(Universität Tübingen)

Additional Contribution

Perfect and Imperfect Nominals

Fritz Hamm

(Universität Tübingen)

Additional Contribution

Coreference and Contextually Restricted Quantification

Jeroen Groenendijk, Martin Stokhof and Frank

Veltman

(Universiteit van Amsterdam)

Additional Contribution

On Context and Identity

Paul Dekker

(Universiteit van Amsterdam)